OpenStudy (anonymous):

Write the equation of the line that passes through the points (–1, –4) and (2, 5).

OpenStudy (anonymous):

find the slope first... in this case the slope is 3... now plug that into Y=mx+b format, to find the y -int.... that would be -1... so the answer is y=3x-1

OpenStudy (2bornot2b):

\[y+4=\frac{-4-5}{-1-2}(x+1)\] Now simplify the expression, and you will get your equation of line